• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name N Toro

    Log entry time 08:27:34 on July 9,2010

    Entry number 331035

    keyword=owl shift summary (abridged)

    Owl shift July 9

    Shift crew:
    SL: Natalia Toro
    TO/DAQ: Neil Goeckner-Wald
    Analysis: Sergey Abrahimyan

    On swing and owl shift we have completed the majority of the optics run plan; we are finishing a few runs during day shift at 1.16 GeV and 1.0454 GeV p0 settings. The latter required additional approval because of beamline-related FSD settings on septum, which are now resolved.

    [Sorry for the brief summary; Rouven will post a list of all optics runs during swing, owl, and day as soon as optics running is completed]